OpenStudy (moonlitfate):

If x= 3/7 is a root of 49x^3 - 189x^2+135x -27 = 0, use synthetic division to factor the polynomial completely and list all real solutions of the equations.

OpenStudy (moonlitfate):

I know that since x = 3/7, that means that (7x-3) is a factor.

OpenStudy (amorfide):

okay so write down (7x-3) okay so 7x times what gives 49x^3?

OpenStudy (moonlitfate):

7x^2

OpenStudy (amorfide):

so the fist value in your second bracket is 7x^2 (7x-3)(7x² now multiply 7x² by -3 then how many more do you need to get 189x²?

OpenStudy (amorfide):

-189x² i meant

OpenStudy (moonlitfate):

Oh, so just use long division. Working out the answer to that part right now.

OpenStudy (anonymous):



OpenStudy (moonlitfate):

I got 7x^2-24x+9

OpenStudy (amorfide):

you are right

OpenStudy (amorfide):

you might be able to factorise the quadratic you have

OpenStudy (moonlitfate):

Factored 7x^2-24x + 9 into: (7x-3) (x-3); so the factors are: (7x-3)^2 (x-3). So the roots are x= 3/7 and x=3. :)

OpenStudy (amorfide):

well done!

OpenStudy (amorfide):

x does not = 7 how did you get that guy?

OpenStudy (anonymous):

Oops, You're correct! 21/7 = 3

OpenStudy (moonlitfate):

Thanks for the help @amorfide :)

OpenStudy (amorfide):

anytime! just happy to help